A 3.0 cm × 3.0 cm × 3.0 cm box with its edges aligned with the xyz-axes is in the electric field E⃗ =(350x+150)i^N/C, where x is in meters.  What is the net electric flux through the box?
Answer #1

here,

volume of box , V = 3 cm * 3 cm * 3cm

area of one face , A = 0.03 ^2 m^2

A = 0.0009 m^2

the net electric flux , Flux = E * A

flux = (350*0.03+150) * A - (350 * 0 + 150)* A

flux = 350 * 0.03 * 0.0009

flux = 9.45 * 10^-3 Nm^2/C

the net electric flux is 9.45 * 10^-3 Nm^2/C